OpenStudy (selena2345):

Ms.Hobson agrees to pay 11% interest on a loan for her new automobile. She is charged $2200 interest on a loan for 1 year. How much did she borrow ?

OpenStudy (ash2326):

She is charged $2200 interest on a loan for 1 year Let she borrowed x , then interest to be paid in a year \[I=x\times 11\%\] interest is 2200 \[2200=x\times \frac{11}{100}\] Can you find x?

OpenStudy (selena2345):

how do i find x?

OpenStudy (ash2326):

First step: Multiply both sides by 100, what would you get?

OpenStudy (selena2345):

which sides

OpenStudy (ash2326):

\[2200\times 100=x\times \frac{11}{100}\times 100\]

OpenStudy (selena2345):

2200 x 100=220000 11/100 x 11=11

OpenStudy (ash2326):

\[220000=x\times 11\] Now divide both sides by 11 to find x

OpenStudy (selena2345):

220000/11=20000 11/11=1

OpenStudy (ash2326):

so what you got for x?

OpenStudy (selena2345):

x=20000

OpenStudy (ash2326):

Correct, this is the amount borrowed. Do you understand?

OpenStudy (selena2345):

yes

OpenStudy (selena2345):

Thank You Ash2326

OpenStudy (ash2326):

Welcome :)